Forecast the demand for the rice for a country for the year 2021 on the basis of 7-year moving average and least square method, data given in table:


Year 2014 2015 2016 2017 2018 2019 2020


Population (millions) 10 12 15 20 25 30 40


Rice consumed (million tonnes) 40 50 60 70 80 90 100



1
Expert's answer
2021-12-23T10:52:28-0500


Mean of x values:

"=(10+12+15+20+25+30+40)\/7=21.7"

Mean of y values:

"=(40+50+60+70+80+90+100)\/7=70"

The normal equations are:

"y=an+b\Sigma x"

"xy=a\Sigma x +b\Sigma x^2"

Subst. into the normal equations:

"7a+152b=490"

"152a+3994b=12000"

"1064a+23104b-(1064a+27958b)=74480-84000"

"b=2."

"7a=490-152(2)"

"a=26.57"

"y=a+bx"

"y=26.57+2x"